### v2

# HRET: A Self-Evolving LLM Evaluation Toolkit for Korean

Abstract
Recent advancements in Korean large language models (LLMs) have spurred numerous benchmarks and evaluation methodologies, yet the lack of a standardized evaluation framework has led to inconsistent results and limited comparability. To address this, we introduce HRET (Haerae Evaluation Toolkit), an open-source, self-evolving evaluation framework tailored specifically for Korean LLMs. HRET unifies diverse evaluation methods, including logit-based scoring, exact-match, language-inconsistency penalization, and LLM-as-a-Judge assessments. Its modular, registry-based architecture integrates major benchmarks (HAE-RAE Bench, KMMLU, KUDGE, HRM8K) and multiple inference backends (LiteLLM, HuggingFace, OpenAI-compatible endpoints). With automated pipelines for continuous evolution, HRET provides a robust foundation for reproducible, fair, and transparent Korean NLP research.
HRET: A Self-Evolving LLM Evaluation Toolkit for Korean
Hanwool Lee1,2  Sooyong Kim1  Dasol Choi1,3  Sangwon Baek1,4 Seunghyeok Hong1,5  Ilgyun Jeong1,6  Naeun Lee1  Inseon Hwang1  Guijin Son1,7 1HAERAE 2Shinhan Securities  3Yonsei University  4Catius  5Hankuk University of Foreign Studies 6Korea University  7OneLineAI gksdnf424@gmail.com
Figure 1: Architecture overview of HRET, showing the registry-based modular design that integrates datasets, models, inference strategies, and evaluation methods through a unified pipeline.
1 Introduction
Mature evaluation practices are essential for accurately tracking advancements in the capabilities of large language models (LLMs). Biased or skewed assessments can mislead the research community and hinder shared goals of technological progress. In this line, the English NLP community has long pursued credible, standardized benchmarks—starting with the GLUE and SuperGLUE leaderboards (Wang et al., 2018, 2019), followed by LM-Eval Harness (Gao et al., 2024), HELM (Liang et al., 2023), and Chatbot Arena—all aimed at providing fair, comparable evaluations under consistent conditions. Now, as Korean NLP research increases in volume, they face similar challenges: Technical reports for foundational models often yield inconsistent metrics, complicating direct comparisons. This underscores the need for a unified evaluation framework tailored to Koreans.
Korean benchmarks vary considerably in their evaluation setups—HAE-RAE Bench utilizes a log-probability-based evaluation, while KMMLU employs exact-match criteria. Such methodological discrepancies often compel researchers and practitioners to customize evaluations, complicating direct comparisons and compromising interpretability. For example, when evaluating identical models like Gemma2-27B on the same benchmarks, different research groups report significantly different results (as shown in Table 1), highlighting the urgent need for standardization.
Additionally, existing evaluations frequently overlook key linguistic nuances unique to Korean, such as intricate morphology, context-dependent word segmentation, and extensive honorific systems. These language-specific challenges require specialized consideration in evaluation methodologies to ensure accurate performance assessment of Korean LLMs.
To address these challenges, we introduce HRET (Haerae Evaluation Toolkit), a standardized, open-source toolkit specifically tailored for evaluating Korean LLMs. HRET makes several key contributions:
  • Unified and reproducible evaluation: HRET standardizes dataset handling, evaluation methods, and benchmarking procedures, ensuring transparency and consistent outcomes.
  • Self-Evolving Mechanism: Its modular registry-based design allows HRET to continuously integrate new benchmarks, evaluation methodologies, and inference backends with minimal effort.
  • Multi-method evaluation: Supports multiple complementary evaluation methods—including logit-based scoring, exact-match evaluation, and LLM-as-a-judge—to comprehensively assess model performance.
  • Ease of Use: Intuitive API and streamlined command-line interfaces simplify evaluation workflows, making it accessible for both researchers and industry practitioners.
By unifying prominent benchmarks and diverse inference backends—including HuggingFace, vLLM, and OpenAI-compatible APIs—HRET significantly improves reproducibility, comparability, and accessibility in Korean NLP research. The toolkit and comprehensive documentation will be publicly released under the Apache License 2.0 upon acceptance to encourage widespread adoption and community contributions.
Report Model Datasets
GSM8K MMLU KMMLU HumanEval MBPP
EXAONE 3.5 Report Research et al. (2024) Qwen2.5-32B 92.00 81.40 62.10 89.00 88.90
Gemma2-27B 84.20 74.80 53.80 79.30 80.70
Kanana Report Bak et al. (2025) Qwen2.5-32B 82.41 83.10 63.15 75.16 73.40
Gemma2-27B 74.37 75.45 51.16 51.22 64.60
Table 1: Performance discrepancies reported for the Gemma-2-27B model between EXAONE 3.5 and Kanana technical reports, highlighting reproducibility challenges in Korean NLP evaluations.
2 Related Work and Background
2.1 LLM Benchmarks and Leaderboards
The English LLM ecosystem has established robust evaluation frameworks like Evalverse Kim et al. (2024b) and deepeval AI (2025). The lm-evaluation-harness Gao et al. (2024) by Upstage has been particularly influential, powering their Open LLM Leaderboard which has accelerated global LLM development. Its success highlights how easy-to-use evaluation tools can democratize model development - lowering barriers for smaller research labs and startups to participate in advancing the field. Similarly, Chinese frameworks like FlagEval Team (2025) have adopted this accessibility-first approach to systematically advance their language-specific ecosystems.
2.2 Reproducibility Challenges
Current Korean NLP evaluations exhibit considerable reproducibility challenges. A notable example is the inconsistent results reported for widely used models such as Gemma2-27B Team et al. (2024) and Qwen2.5-32B Yang et al. (2024) across various technical reports from major Korean organizations. As illustrated in Table 1, evaluations of identical benchmarks using these same models yield significantly different outcomes depending on the evaluating institution. For instance, the EXAONE 3.5 report Research et al. (2024) and Kanana report Bak et al. (2025) report performance discrepancies ranging from around 1–10 percentage points across multiple tasks, including GSM8K, KMMLU, HumanEval, and MBPP.
These discrepancies stem from variations in inference settings, prompt engineering, and evaluation implementations. The situation is further complicated by Korean’s linguistic features—complex morphology, context-dependent segmentation, and honorific systems—which are handled inconsistently across evaluations. Existing benchmarks lack standardized guidelines addressing these Korean-specific challenges, resulting in fragmented evaluation practices that impede meaningful progress in Korean NLP research.
Feature lm-eval-harness Biderman et al. (2024) OpenAI Evals Liu et al. (2024) Evalverse Kim et al. (2024b) DeepEval AI (2025) HRET (Ours)
Primary Focus Academic benchmarks OpenAI model evaluation Unified evaluation hub Unit/regression testing Korean LLM evaluation
Developer EleutherAI OpenAI UpstageAI Confident AI HAERAE
Architecture Task/Metric modules Evaluation scripts Submodule integration Test case-based Registry-based plugins
Built-in Benchmarks 60+ general benchmarks Several public evals Integrated external benchmarks Major benchmarks Korean-focused benchmarks
Evaluation Metrics Traditional metrics Task-specific scripts Integrated metrics 14+ diverse metrics Traditional + LLM-as-judge
User Interface CLI only CLI + Dashboard Slack bot + Web CLI + Web dashboard CLI + API
Automation Script-based API-based No-code automation CI/CD integration Pipeline Runner
Unique Strength High reproducibility OpenAI optimization Multi-tool integration Testing-focused Registry-based extensibility
Table 2: Feature comparison of HRET with major existing LLM evaluation frameworks, highlighting its registry-based extensibility and Korean-specific benchmarks.
2.3 Comparison with Existing Systems
To position HRET within the landscape of LLM evaluation frameworks, Table 2 compares its features with other systems. This comparison highlights HRET’s unique contributions to the field, particularly its focus on Korean language evaluation and its registry-based modular architecture.
HRET differentiates itself through its specific focus on Korean language evaluation challenges, its registry-based architecture enabling easy extension, and its integrated pipeline approach that combines dataset handling, model inference, scaling, and evaluation into a unified framework.
3 HRET Framework
HRET (Haerae Evaluation Toolkit) addresses critical reproducibility and comparability challenges faced in evaluating Korean large language models (LLMs). By providing a standardized yet flexible evaluation environment, HRET fosters consistency, transparency, and accessibility in Korean NLP research. Figure 1 illustrates its modular and registry-based architecture, seamlessly integrating datasets, evaluation methods, inference strategies, and model backends into a coherent evaluation pipeline.
HRET is explicitly designed for: (1) researchers who require standardized evaluation methods for benchmarking Korean language models; (2) developers looking for robust tools to streamline model assessment and selection; and (3) industry practitioners aiming to reliably deploy Korean LLMs in production.
3.1 Novelty and Design Principles
HRET introduces several innovative approaches to address the challenges of Korean LLM evaluation:
  • Self-Evolving Mechanism: HRET continuously adapts by automatically discovering and integrating new evaluation methods, datasets, and inference backends at runtime through (1) Automated Component Discovery for runtime integration without core code modification, (2) Dynamic Pipeline Construction that adapts to new scenarios without predefined workflows, and (3) Continuous Integration Workflows that validate components against established benchmarks.
  • Registry-Based Modularity: Core components—including dataset loaders, evaluation methods, scaling strategies, and inference backends—are individually modularized and managed through registries, allowing researchers to effortlessly extend the toolkit’s capabilities with minimal code changes.
  • Concurrency and Scalability: Leveraging asynchronous HTTP requests (via httpx) and multi-threading, HRET significantly enhances throughput during large-scale evaluations, essential for computationally intensive Korean benchmarks.
  • Unified API: HRET simplifies interaction through consistent Evaluator and PipelineRunner interfaces, reducing the entry barrier for both novice researchers and industry practitioners.
  • Korean Focus: The framework specifically targets Korean language benchmarks and evaluation needs, with built-in support for major Korean evaluation datasets and specialized evaluation methods like language consistency checking.
Category Benchmark Description
General HAE-RAE Son et al. (2023) Evaluation of Korean language understanding, focusing on vocabulary knowledge, reading comprehension, and general knowledge across multiple domains.
KMMLU Son et al. (2024b) Assessment of multi-domain knowledge across 45 categories including STEM, humanities, applied sciences, and other fields in Korean language context.
CLIcK Kim et al. (2024a) Measurement of cultural and linguistic proficiency through 1,995 QA pairs from Korean exams and textbooks, organized into eleven distinct categories.
Dialogue KUDGE Son et al. (2024a) Evaluation of dialogue generation capabilities, with emphasis on coherence and contextual appropriateness in Korean conversational settings.
Specialized K2-Eval Son et al. (2024a) Assessment of domain-specific knowledge in professional, technical, and academic fields requiring specialized Korean terminology.
HRM8K Ko et al. (2025) Evaluation of mathematical reasoning using 8,011 problems, combining 1,428 native Korean math exam questions and 6,583 translated problems from other benchmarks.
Table 3: Descriptions of benchmarks supported in HRET.
3.2 System Architecture and Components
HRET provides a unified interface for evaluating major Korean NLP benchmarks (detailed in Table 3), following a standardized five-stage workflow:
  1. 1.
    Dataset Loading: The Dataset Registry unifies Korean NLP datasets with consistent formatting, supporting standard benchmarks and offering a flexible generic_file loader for custom datasets without requiring dedicated classes.
  2. 2.
    Model Loading: The Backend Registry connects to various backends (HuggingFace, OpenAI-compatible, LiteLLM), providing a unified interface that abstracts provider-specific details while supporting generation, judge, and reward functionalities.
  3. 3.
    Inference with Scaling: Scaling Techniques enhance evaluation quality through (1) Beam Search  (Chen et al., 2024) for exploring different decoding paths, (2) Best-of-N Sampling  (Stiennon et al., 2022) for selecting optimal outputs, and (3) Majority Voting  (Wang et al., 2023) for enhancing stability through consensus.
  4. 4.
    Evaluation: Generated predictions are assessed using multiple approaches: (1) String-matching for reference comparison, (2) Logit-based Evaluation for confidence analysis, and (3) LLM-as-a-Judge for open-ended assessment. Additional Evaluation Enhancements include Language penalizer for detecting non-Korean responses and Answer extraction for identifying final answers from reasoning chains.
  5. 5.
    Results Reporting: Runtime Components orchestrate the pipeline, with the Runner combining components from all registries and the Evaluator implementing a facade pattern providing a simplified API.
This architecture ensures consistency across evaluations while providing flexibility through pluggable components at each stage.
3.3 Advanced Features and Usage
HRET supports multiple inference backends through a consistent interface:
  • HuggingFace Transformers: Provides direct access to thousands of models, including Korean-specific models like KoGPT and HyperCLOVA.
  • vLLM: Optimized for high-throughput evaluation scenarios with large Korean benchmarks.
  • OpenAI-compatible endpoints: Ensures compatibility with API-based services from major Korean AI providers.
  • LiteLLM: Offers unified access to multiple model providers, simplifying cross-vendor comparisons.
Beyond the standard Evaluator API, HRET provides a flexible PipelineRunner for advanced customization of evaluation workflows. This interface enables:
  • Multi-stage evaluations where outputs from one model serve as inputs to another
  • Comprehensive model comparisons across different benchmarks and metrics
  • Custom workflows with specialized processing steps for Korean-specific evaluations
  • Ensemble evaluations combining complementary models or methodologies
3.4 Usage Examples
To illustrate the ease of use and practical value of HRET, we provide a brief usage example evaluating a Korean LLM on the HAE-RAE Bench benchmark using our command-line interface (CLI):
python -m llm_eval.evaluator \
--model huggingface \
--model_params ’{"model_name_or_path": "google/gemma-3-1b-it"}’ \
--dataset haerae_bench \
--subset csat_geo \
--split test \
--evaluation_method string_match \
--output_file results.json
This command: (1) Loads the specified benchmark dataset (‘HAE-RAE Bench‘, subset ‘csat_geo‘), (2) Instantiates the chosen HuggingFace model (‘Gemma-3-1b-it‘), (3) Evaluates the model outputs using the exact-match criteria, and (4) Outputs evaluation results in JSON format.
Additional detailed usage scenarios and advanced functionalities (Python API, multi-method evaluations, scaling methods) are provided in Appendix A.
4 System Validation and Roadmap
4.1 QA Testing and Benchmarking
HRET underwent rigorous validation to ensure reliability and usability:
QA Testing
We conducted extensive functional testing through a rigorous QA process, ensuring all components performed reliably under diverse configurations and use cases.
Real-World Deployment
The framework was deployed as the backend system for the HAE-RAE Reasoning Challenge leaderboard, demonstrating its robustness in a production environment with various model submissions and evaluation criteria.
Performance Benchmarking
We conducted extensive performance testing to ensure the system maintains efficiency with large-scale Korean benchmark datasets, validating that the asynchronous and multi-threading capabilities deliver the expected throughput improvements.
4.2 Current Status and Roadmap
HRET is under active development, currently supporting all benchmarks mentioned and the evaluation methodologies described. The ongoing roadmap includes strengthening the self-evolving mechanism through enhanced implementation of the hret-agent, which leverages the plugin-based registry architecture for improved scalability. Additionally, HRET aims to deepen integration and interoperability with other open-source LLM libraries, promoting broader adoption and community collaboration.
5 Experiments
We conduct two experiments to highlight how HRET effectively addresses critical challenges specific to Korean LLM evaluations.
5.1 Language Consistency Evaluation with Language Penalizer
Multilingual models often respond in non-Korean languages, particularly English, skewing evaluations on Korean tasks. HRET’s language penalizer detects and penalizes such responses, ensuring results accurately reflect Korean language capabilities. Figure 4 shows significant performance adjustments for several models on the K2-Eval benchmark when language penalization is enabled.
Figure 2: Performance comparison with and without language penalizer on K2-Eval. Qwen models show reductions when enforced, while EXAONE and Llama models maintain consistent scores.
5.2 Multi-Method Evaluation Approaches
HRET supports multiple evaluation methods tailored for different task types. Table 4 compares the performance of EXAONE-3.5-2.4B across evaluation methods on HRM8K and CLIcK benchmarks.
Benchmark String-Match Math-Verify
HRM8K 2.21 14.11
CLIcK 27.31 28.67
Table 4: Comparison of evaluation methods for EXAONE-3.5-2.4B across benchmarks.
Key insights include:
  • Performance varies significantly with evaluation method, particularly for mathematical reasoning tasks (HRM8K shows a 6.4x improvement with Math-Verify).
  • For language understanding datasets like CLIcK, evaluation methods yield more consistent results (difference of only 1.36%).
  • Task-appropriate evaluation metrics provide more accurate assessments of model capabilities.
  • HRET enables researchers to select the most suitable evaluation method for different Korean language tasks.
These experiments demonstrate how HRET’s diverse evaluation methodologies reveal important performance nuances that might be missed with a single evaluation approach.
6 Conclusion
We introduce HRET, a self-evolving, modular evaluation toolkit tailored specifically for Korean large language models. HRET standardizes methodologies, improves reproducibility, and addresses linguistic nuances specific to Korean. Future work will include expanding benchmarks for domain-specific linguistic analyses, deeper integration with emerging LLM libraries, and enhancing visualization tools to facilitate detailed error diagnostics. Ultimately, HRET aims to establish a robust foundation for Korean NLP research, supporting sustained progress and collaboration within the community.
Limitations
While HRET provides a standardized evaluation framework for Korean LLMs, several limitations remain:
First, despite integrating major Korean benchmarks, coverage in specialized domains (legal, medical, scientific) remains limited. Second, the language penalizer may struggle with mixed-language outputs or code-switching common in technical contexts. Third, computational requirements for scaling techniques can be substantial, potentially limiting accessibility.
Current evaluation metrics, particularly string matching, may not fully capture Korean’s rich morphological structures. Additionally, as new Korean LLMs and benchmarks emerge, maintaining compatibility requires continuous updates. Finally, the toolkit offers limited tools for analyzing model errors specific to Korean linguistic phenomena.
We are addressing these limitations through ongoing development and welcome community contributions to enhance HRET’s capabilities.
Appendix A Usage Examples and Tutorials
This section provides advanced examples and tutorials for using HRET beyond the basic functionality introduced in the main paper.
A.1 Advanced Backend Usage
HRET supports multiple backend integrations beyond standard HuggingFace models:
A.1.1 LiteLLM and OpenAI-Compatible APIs
Access over 30 LLM providers and self-hosted servers:
# Using LiteLLM for cloud-based models
results = evaluator.run(
model="litellm",
model_params={
"model_name": "gpt-4",
"api_key": "your-api-key-here",
dataset="haerae_bench",
evaluation_method="string_match"
# Using OpenAI-compatible endpoints for self-hosted models
results = evaluator.run(
model="openai",
model_params={
"model_name": "local-llama-2-13b",
"api_base": "http://localhost:8000/v1",
dataset="haerae_bench",
evaluation_method="string_match"
A.2 Evaluation Methods
A.2.1 String Match and Log Probability Evaluation
HRET offers flexible text matching with various normalization options:
# String matching with normalization options
results = evaluator.run(
model="huggingface",
model_params={"model_name_or_path":"Qwen/Qwen2.5-3B-Instruct"},
dataset="haerae_bench",
evaluation_method=’string_match’,
evaluator_params={
"ignore_case": True,
"ignore_punctuation": True,
"extract_answer": True,
# Log probability evaluation for MCQA
results = evaluator.run(
model="huggingface",
dataset="haerae_bench",
evaluation_method=’log_likelihood’,
dataset_params={"base_prompt_template": "{query} ### Answer:"},
A.2.2 Math Verification
Semantic equivalence checking for mathematical expressions:
results = evaluator.run(
model="huggingface",
model_params={"cot": True}, # Enable Chain-of-Thought
dataset="hrm8k",
evaluation_method="math_verify",
evaluator_params={"extract_final_answer": True}
A.3 Advanced Features
A.3.1 Scaling Techniques and Chain-of-Thought
Enhance evaluation with scaling methods and reasoning:
# Self-consistency evaluation
results = evaluator.run(
model="huggingface",
dataset="haerae_bench",
scaling_method=’self_consistency’,
scaling_params={"num_samples": 5}
# CoT with specific trigger phrase
results = evaluator.run(
model="huggingface",
dataset="haerae_bench",
model_params={
"cot": True,
"cot_trigger": "Let’s think step by step."
A.3.2 Model Comparison and Korean Language Support
Compare multiple models and handle language-specific evaluation:
# Define models to compare
models = [
{"model": "huggingface", "model_params": {"model_name_or_path": "Qwen/Qwen2.5-7B-Instruct"}, "name": "Qwen-7B"},
{"model": "huggingface", "model_params": {"model_name_or_path": "kakaocorp/kanana-nano-2.1b-instruct"}, "name": "Kanana-2.1B"}
# Run comparison
comparison_results = evaluator.compare(
models=models,
dataset="haerae_bench",
evaluation_method="string_match"
# Korean language penalizer
results = evaluator.run(
model="huggingface",
dataset="k2_eval",
evaluation_method="string_match",
evaluator_params={
"language_penalizer": True,
"target_language": "ko"
These examples demonstrate HRET’s capabilities for Korean language model evaluation beyond the basic functionality presented in the main paper.
These examples demonstrate HRET’s capabilities for Korean language model evaluation beyond the basic functionality presented in the main paper. Upon publication, comprehensive tutorials and documentation in both Korean and English will be made publicly available in the project repository, including additional examples, best practices, and implementation guidelines.
Appendix B Author Contributions
Hanwool Lee led the project, system design, core development, and quality assurance.
Sooyong Kim contributed extensively to quality assurance, dataset implementation, and scaling methods.
Dasol Choi contributed to paper organization and writing, as well as dataset and evaluation method implementation.
Sangwon Baek worked on backend implementation, tutorial development, and quality assurance.
Seunghyeok Hong contributed to tutorial organization and writing, and implemented several datasets.
Ilgyun Jung developed the LLM-as-a-Judge module and implemented multiple datasets.
Inseon Hwang contributed to dataset implementation and CI/CD pipeline integration.
Naeun Lee worked on backend implementation.
Guijin Son organized the team, contributed to dataset implementation and paper writing, and led the planning and specification of LLM evaluation features.
All authors participated in discussions, testing, and refining the HRET framework.