### v3

# Redefining Evaluation Standards: A Unified Framework for Evaluating the Korean Capabilities of Language Models

Abstract
Recent advancements in Korean large language models (LLMs) have driven numerous benchmarks and evaluation methods, yet inconsistent protocols cause up to 10 pp performance gaps across institutions. Overcoming these reproducibility gaps does not mean enforcing a one-size-fits-all evaluation. Rather, effective benchmarking requires diverse experimental approaches and a framework robust enough to support them. To this end, we introduce HRET (Haerae Evaluation Toolkit), an open-source, registry-based framework that unifies Korean LLM assessment. HRET integrates major Korean benchmarks, multiple inference backends, and multi-method evaluation, with language consistency enforcement to ensure genuine Korean outputs. Its modular registry design also enables rapid incorporation of new datasets, methods, and backends, ensuring the toolkit adapts to evolving research needs. Beyond standard accuracy metrics, HRET incorporates Korean-focused output analyses—morphology-aware Type–Token Ratio (TTR) for evaluating lexical diversity and systematic keyword-omission detection for identifying missing concepts—to provide diagnostic insights into language-specific behaviors. These targeted analyses help researchers pinpoint morphological and semantic shortcomings in model outputs, guiding focused improvements in Korean LLM development.
Redefining Evaluation Standards: A Unified Framework for Evaluating the Korean Capabilities of Language Models
Hanwool Lee1,2thanks: Equal contribution., Dasol Choi1,311footnotemark: 1, Sooyong Kim1, Ilgyun Jung1,4, Sangwon Baek1,6, Guijin Son1,7, Inseon Hwang1, Naeun Lee1, Seunghyeok Hong1,5thanks: Corresponding author. 1MODULABS, HAERAE LAB 2Shinhan Securities  3Yonsei University  4Catius  5Hankuk University of Foreign Studies 6Korea University  7OneLineAI gksdnf424@gmail.com
Figure 1: Architecture overview of HRET, showing the registry-based modular design that integrates datasets, models, inference strategies, and evaluation methods through a unified pipeline.
1 Introduction
Recent advances in large language models (LLMs) demand rigorous, reproducible evaluation, yet real-world benchmarking often diverges from official protocols Liang et al. (2022); Gao et al. (2024). In Korean NLP, identical models have exhibited performance gaps up to 10 percentage points across institutions Research et al. (2024); Bak et al. (2025), driven by variations in prompt templates, inference settings, and evaluation criteria. We argue that true reproducibility arises not from a rigid framework but from a toolkit that flexibly supports multiple evaluation paradigms—few-shot versus log-likelihood scoring, exact-match versus generative judgment, and test-time scaling techniques—so that assessments can adapt to both model characteristics and deployment contexts.
Consider KMMLU Son et al. (2024b), whose official log-likelihood/direct-generation implementation in lm-eval-harness is rarely used unchanged: practitioners often augment it with n-shot in-context examples Brown et al. (2020), inject chain-of-thought prompts, or apply self-consistency Wang et al. (2023) to better capture model reasoning. Such customization highlights the need for an extensible evaluation framework that integrates diverse Korean benchmarks and inference backends while exposing configurable strategies for prompt design, scoring, and aggregation.
To meet this need, we present HRET (Haerae Evaluation Toolkit), an open-source, registry-based framework for unified Korean LLM assessment. HRET standardizes data ingestion and result reporting, supports string matching, partial matching, log-likelihood, mathematical verification, and LLM-as-a-judge methods Zheng et al. (2023), and applies a language-penalty mechanism to discourage non-Korean outputs. Importantly, it embeds Korean-specific diagnostics—morphology-aware Type–Token Ratio for lexical diversity and systematic keyword-omission detection—to surface linguistic phenomena unique to Korean. Its modular registry design enables rapid onboarding of new datasets, methods, and backends, future-proofing evaluation as the field evolves.
We validate HRET’s robustness and efficiency across diverse hardware configurations, demonstrating substantial reductions in evaluation variability and improved throughput. HRET thus offers a practical solution for reproducible, adaptable benchmarking in Korean LLM research.
Report Model Datasets
GSM8K MMLU KMMLU HumanEval MBPP
EXAONE 3.5 Report Research et al. (2024) Qwen2.5-32B 92.00 81.40 62.10 89.00 88.90
Gemma2-27B 84.20 74.80 53.80 79.30 80.70
Kanana Report Bak et al. (2025) Qwen2.5-32B 82.41 83.10 63.15 75.16 73.40
Gemma2-27B 74.37 75.45 51.16 51.22 64.60
Table 1: Performance discrepancies for identical models Qwen et al. (2025); Team et al. (2024) between EXAONE 3.5 and Kanana technical reports, highlighting reproducibility challenges in Korean NLP evaluations.
2 Related Work and Background
2.1 LLM Benchmarks and Leaderboards
The English LLM ecosystem has established robust evaluation frameworks like Evalverse Kim et al. (2024b) and deepeval AI (2025). The lm-evaluation-harness Gao et al. (2024) by Upstage has been particularly influential, powering their Open LLM Leaderboard which has accelerated global LLM development. Its success highlights how easy-to-use evaluation tools can democratize model development - lowering barriers for smaller research labs and startups to participate in advancing the field. Similarly, Chinese frameworks like FlagEval Team (2025) have adopted this accessibility-first approach to systematically advance their language-specific ecosystems.
2.2 Reproducibility Challenges
Evaluations of Korean language models often show inconsistent results across institutions, with the same models reporting performance gaps of up to 10 percentage points. For instance, Gemma2-27B and Qwen2.5-32B Qwen et al. (2025) achieve scores that differ by 1–10 pp on GSM8K, KMMLU, HumanEval and MBPP when comparing the EXAONE 3.5 report Research et al. (2024) and the Kanana report Bak et al. (2025) (Table 1). This variability underscores the absence of consistent evaluation practices tailored to Korean LLMs, complicating direct comparisons and hindering reproducibility.
These gaps arise from varying inference settings, prompt designs, and evaluation implementations. On top of that, Korean’s own characteristics—its rich morphology, context-dependent tokenization, and honorific system—make consistent handling difficult without standardized, Korean-focused guidelines. Together, these factors exacerbate result variability and impede reliable benchmarking across research groups.
Feature lm-eval-harness Biderman et al. (2024) OpenAI Evals Liu et al. (2024) Evalverse Kim et al. (2024b) DeepEval AI (2025) HRET (Ours)
Primary Focus Academic benchmarks OpenAI model evaluation Unified evaluation hub Unit/regression testing Korean LLM evaluation
Developer EleutherAI OpenAI Upstage Confident AI HAERAE
Architecture Task/Metric modules Evaluation scripts Submodule integration Test case-based Registry-based plugins
Built-in Benchmarks 60+ general benchmarks Several public evals Integrated external benchmarks Major benchmarks Korean-focused benchmarks
Evaluation Metrics Traditional metrics Task-specific scripts Integrated metrics 5+ diverse metrics 5+ diverse metrics
User Interface CLI only CLI + Dashboard Slack bot + Web CLI + Web dashboard CLI + API
Automation Script-based API-based No-code automation CI/CD integration API-based, CI/CD integration
Unique Strength High reproducibility OpenAI optimization Multi-tool integration Testing-focused Korean-specific + Registry extensibility
Table 2: Feature comparison of HRET with major existing LLM evaluation frameworks, highlighting its registry-based extensibility and Korean-specific benchmarks.
2.3 Comparison with Existing Systems
To position HRET within the landscape of LLM evaluation frameworks, Table 2 compares its features with other systems. This comparison highlights HRET’s unique contributions to the field, particularly its focus on Korean language evaluation and its registry-based modular architecture.
HRET differentiates itself through three key aspects: (1) Korean-specific capabilities including language consistency enforcement and concise morphological analysis–based output diagnostics, (2) registry-based extensibility that enables rapid component integration compared to traditional monolithic architectures, and (3) unified evaluation pipeline that unifies fragmented Korean LLM assessment practices. Unlike existing frameworks that focus on general-purpose evaluation, HRET addresses specific reproducibility challenges observed in Korean NLP research (Table 1).
3 HRET Framework
HRET (Haerae Evaluation Toolkit) addresses critical reproducibility and comparability challenges faced in evaluating Korean large language models (LLMs). By providing a standardized yet flexible evaluation environment, HRET fosters consistency, transparency, and accessibility in Korean NLP research. Figure 1 illustrates its modular and registry-based architecture, seamlessly integrating datasets, evaluation methods, inference strategies, and model backends into a coherent evaluation pipeline.
HRET is explicitly designed for: (1) researchers who require standardized evaluation methods for benchmarking Korean language models; (2) developers looking for robust tools to streamline model assessment and selection; and (3) industry practitioners aiming to reliably deploy Korean LLMs in production.
3.1 Design Principles
HRET introduces several innovative approaches to address the challenges of Korean LLM evaluation:
  • Registry-Based Modularity: Core components—including dataset loaders, evaluation methods, scaling strategies, and inference backends—are individually modularized and managed through registries, allowing researchers to effortlessly extend the toolkit’s capabilities with minimal code changes compared to traditional monolithic architectures.
  • Unified Standardization: HRET systematically standardizes fragmented Korean LLM evaluation practices by providing consistent interfaces, reproducible workflows, and transparent benchmarking procedures across different research environments.
  • Korean Language Consistency: Penalizes any response not in Korean by applying a score penalty, ensuring evaluation focuses strictly on Korean outputs.
  • Concurrency and Scalability: Leveraging asynchronous HTTP requests (via httpx) and multi-threading, HRET significantly enhances throughput during large-scale evaluations, essential for computationally intensive Korean benchmarks.
  • Unified API: HRET simplifies interaction through consistent Evaluator and PipelineRunner interfaces, reducing the entry barrier for both novice researchers and industry practitioners.
3.2 System Architecture and Components
HRET provides a unified interface for evaluating major Korean NLP benchmarks including HAE-RAE Bench, KMMLU, KUDGE, HRM8K, CLIcK, and others (see Appendix A for complete list), following a standardized five-stage workflow:
  1. 1.
    Dataset Loading: The Dataset Registry unifies Korean NLP datasets with consistent formatting, supporting standard benchmarks and offering a flexible generic_file loader for custom datasets without dedicated classes.
  2. 2.
    Model Loading: The Backend Registry connects to various backends (HuggingFace, vLLM, OpenAI-compatible, etc), providing a unified interface that abstracts provider-specific details while supporting generation, judge, and reward functionalities.
  3. 3.
    Inference with Scaling: Scaling Techniques enhance evaluation quality through (1) Beam Search for exploring different decoding paths, (2) Best-of-N Sampling for selecting optimal outputs, and (3) Majority Voting for enhancing stability through consensus.
  4. 4.
    Evaluation: Generated predictions are assessed using multiple approaches: (1) String-matching, (2) Logit-based Evaluation (3) Mathematical Verification for semantic equivalence in mathematical reasoning tasks, and (4) LLM-as-a-Judge. Additional enhancements include Language consistency enforcement for detecting non-Korean responses and Answer extraction for identifying final answers from reasoning chains.
  5. 5.
    Results Reporting: Runtime Components orchestrate the pipeline, with the Runner combining components from all registries and the Evaluator implementing a facade pattern providing a simplified API.
This architecture ensures consistency across evaluations while providing flexibility through pluggable components at each stage.
3.3 Advanced Evaluation Pipeline
Beyond the standard Evaluator API, HRET provides a flexible PipelineRunner for advanced customization of evaluation workflows. This interface enables multi-stage evaluations where outputs from one model serve as inputs to another, comprehensive model comparisons across different benchmarks and metrics, custom workflows with specialized processing steps for Korean-specific evaluations, and ensemble evaluations combining complementary models or methodologies.
3.4 Korean-Specific Analysis Capabilities
HRET extends beyond standard metrics by integrating a Korean-specific morphological analyzer to deliver four core analyses: (1) Overall Performance with detailed accuracy, sample counts, and prediction distributions; (2) Subset Analysis highlighting variations across task types (e.g., reading comprehension vs. cultural knowledge); (3) Lexical Diversity via Type–Token Ratio (TTR) comparisons between correct and incorrect outputs; and (4) Error Pattern Detection through keyword-based inference of frequent failure modes.
Together, these tools form a diagnostic framework tailored to Korean linguistic characteristics—highlighting morphological limitations and vocabulary gaps that standard accuracy alone cannot capture, and guiding targeted improvements in model training and evaluation.
3.5 Usage Examples
HRET provides intuitive command-line and Python API interfaces. A simple CLI example for evaluating a Korean LLM on HAE-RAE Bench:
python -m llm_eval.evaluator \
--model huggingface \
--model_params ’{"model_name_or_path": "google/gemma-3-1b-it"}’ \
--dataset haerae_bench \
--subset csat_geo \
--split test \
--evaluation_method string_match \
--output_file results.json
This command loads the specified dataset and model, performs string-match evaluation, and outputs results in JSON format. Additional usage scenarios and advanced functionalities are provided in Appendix D.
4 System Evaluation
4.1 Reproducibility Validation
To validate HRET’s ability to address reproducibility challenges in Korean LLM evaluation, we conducted cross-institutional experiments across different hardware configurations and backend implementations. We evaluated three representative models (Qwen2.5-3B Qwen et al. (2025), Qwen3-8B Yang et al. (2025), Llama-3.2-3B-Instruct Meta AI (2024)) on Korean benchmarks using log-likelihood evaluation across four hardware setups with varying CPU architectures, GPU types, and memory configurations.
Model Max Variation Hardware Configs
Qwen2.5-3B 0.0016 (0.56%) 4 setups
Qwen3-8B 0.0028 (1.06%) 4 setups
Llama-3.2-3B 0.0015 (0.62%) 4 setups
Table 3: Hardware robustness summary across environments (detailed table in Appendix C).
Results demonstrate remarkable consistency despite significant hardware variations. Accuracy scores remained stable across different CPU architectures, GPU types, and batch sizes (1–64), with standard deviations under 0.003. Cross-institutional experiments showed identical or near-identical scores, validating HRET’s ability to eliminate evaluation discrepancies observed in Table 1.
Hardware Configuration HAE-RAE CLIcK
AMD EPYC 74F3 + H100 0.1806 0.1343
Intel Xeon + A100-80GB 0.1788 0.1298
Intel Xeon + H200 0.1802 0.1308
Standard Deviation 0.0009 0.0023
Table 4: Backend robustness using vLLM for Qwen2.5-3B.
Backend comparison between HuggingFace and vLLM shows that while absolute scores differ due to implementation differences, results remain highly consistent within each backend ($\sigma<0.003$). Compared to the 1–10 percentage point discrepancies reported between institutions, HRET achieves hardware robustness with standard deviations under 0.003 and cross-institutional consistency when using standardized procedures.
4.2 Multi-Method Evaluation Analysis
Korean language tasks exhibit diverse characteristics that benefit from different evaluation approaches. HRET’s multi-method evaluation framework enables researchers to select the most appropriate assessment strategy for each task type, leading to more accurate performance measurements.
Table 5 compares evaluation methods across different Korean benchmarks using EXAONE-3.5-2.4B Research et al. (2024) as a case study. The results demonstrate significant performance variations depending on the evaluation method employed, with notable differences between task types.
Benchmark String-Match Math-Verify LLM-as-judge(w.gpt-4o-2024-11-20)
HRM8K 2.21 14.11 17.60
CLIcK 27.31 28.67 27.76
Table 5: Comparison of evaluation methods for EXAONE-3.5-2.4B across Korean benchmarks.
Task-Dependent Sensitivity
Mathematical reasoning (HRM8K) sees dramatic score increases—from 2.21% (String-Match) to 17.60% (LLM-as-judge)—highlighting the need for semantic evaluation over exact matching.
Evaluation Robustness
Cultural and linguistic tasks (CLIcK) remain consistent (27.31–28.67%) across methods, indicating lower sensitivity to metric choice.
Methodological Implications
By matching metric to task—semantic methods for reasoning, exact-match for comprehension—HRET balances accuracy and efficiency across diverse Korean language evaluations.“‘
4.3 Test-Time Scaling Techniques
HRET incorporates advanced test-time scaling techniques to enhance evaluation quality through multiple inference strategies. These methods are particularly valuable for Korean language evaluation, where single predictions may not fully capture model capabilities due to linguistic complexity and cultural contexts.
The framework supports three primary approaches: Best-of-N Sampling, Self-Consistency with majority voting, and Beam Search for optimal decoding paths. These techniques can be combined with various evaluation methods, including LLM-as-a-judge assessments.
Table 6 demonstrates test-time scaling effectiveness on HAERAE Reasoning Challenge (HRC), evaluating Korean reasoning across science, mathematics, and quiz domains.
Model Benchmark n=1 n=3 n=5
EXAONE-3.5-2.4B HRC 3.37 5.04 9.24
Qwen2.5-3B-Instruct 11.76 12.60 13.44
Table 6: Test-time scaling results with LLM-as-a-judge(w. gpt-4o-2024-11-20) on HRC.
Results show substantial performance gains through increased sampling, with EXAONE-3.5-2.4B achieving 2.7x improvement at n=5 versus single-shot evaluation. This validates test-time scaling effectiveness for complex Korean reasoning tasks requiring nuanced understanding of context and cultural references.
4.4 Extensibility
Our registry-based design uses simple Python decorators to enable seamless extension. New Models inherit from BaseModel and add @register_model; new Datasets inherit from BaseDataset and add @register_dataset; new Evaluators inherit from BaseEvaluator and add @register_evaluator; and new Test-time scaling methods inherit from BaseScalingMethod and add @register_scaling_method. All decorated classes are auto-registered at runtime, allowing users to plug in new components with zero additional configuration.
To foster community-driven development, HRET actively welcomes external contributions and is committed to continuous maintenance and expansion. This collaborative approach ensures the framework evolves with the Korean LLM landscape while maintaining backward compatibility for existing users.
5 Conclusion
We introduce HRET, a unified evaluation framework that addresses critical reproducibility challenges in Korean LLM evaluation. Through its registry-based architecture and standardized protocols, HRET significantly improves evaluation consistency while supporting diverse assessment strategies. HRET integrates major Korean benchmarks with multiple inference backends and evaluation methods, enabling researchers to select task-appropriate strategies while maintaining reproducibility. The framework’s extensible design and community-driven development ensure it evolves with the Korean LLM landscape, providing a sustainable foundation for reproducible assessment.
Limitations
While HRET provides a standardized evaluation framework for Korean LLMs, several limitations remain. Despite integrating major Korean benchmarks, coverage in specialized domains Hwang et al. (2025) remains limited. The language consistency enforcement may struggle with mixed-language outputs or code-switching common in technical contexts, and computational requirements for test-time scaling techniques can be substantial, potentially limiting accessibility for resource-constrained environments. Current evaluation metrics, particularly string matching, may not fully capture Korean’s rich morphological variations and honorific systems, and the framework’s effectiveness depends on the quality of integrated benchmarks. We are addressing these limitations through ongoing development and welcome community contributions to enhance HRET’s capabilities.
Appendix A Supported Benchmarks
HRET provides unified evaluation support for major Korean language benchmarks spanning diverse domains and task types. The framework’s registry-based architecture enables seamless integration of these benchmarks with consistent preprocessing, evaluation protocols, and result formatting. Table 7 provides detailed descriptions of currently supported benchmarks, organized by evaluation focus.
The benchmark collection covers three primary categories: General benchmarks assess broad Korean language understanding and cultural knowledge, Dialogue benchmarks evaluate conversational capabilities in Korean contexts, and Specialized benchmarks target domain-specific competencies including mathematical reasoning and professional knowledge. This comprehensive coverage enables researchers to conduct thorough evaluations across the spectrum of Korean language model capabilities.
Metric HRET lm-eval-harness
HAE-RAE Bench
Execution Time (sec) 114.93 142.44
Max RAM (MB) 80.96 (<1%) 54.26 (<1%)
Max CPU (%) 525.84 100.0
KMMLU
Execution Time (sec) 303.93 221.04
Max RAM (MB) 104.81 (<1%) 78.76 (<1%)
Max CPU (%) 458.90 199.0
Table 8: Performance comparison on AMD EPYC 74F3 64-Core + NVIDIA H100 with 193GB RAM. Values averaged across 3 runs.
Appendix B Performance Comparison with Existing Tools
Table 8 compares HRET’s computational performance with lm-evaluation-harness using Qwen2.5-3B on identical hardware configurations.
HRET shows mixed performance: faster on HAE-RAE Bench but slower on KMMLU. Higher RAM usage stems from loading datasets as editable objects to enable real-time filtering and experimental modifications through our system. Higher CPU utilization (up to 525%) reflects multi-threading optimization. Both frameworks use <1% of total system memory, indicating efficient resource utilization despite HRET’s flexibility-oriented design.
Appendix C Detailed Reproducibility Results
To complement the summary presented in Section 4.1, we provide the full hardware-wise evaluation results across multiple configurations. This table reports evaluation scores for three representative models using the HuggingFace backend under diverse CPU and GPU settings. These include server-grade (A100, H100) and workstation-grade (RTX A6000, RTX 3090) environments, reflecting real-world variance across institutions.
Model Hardware Setup HAE-RAE CLIcK Batch Size
Qwen2.5-3B AMD EPYC 74F3 + A100×2 0.2847 0.3388 64
AMD EPYC 9354 + RTX A6000 0.2847 0.3388 2
Intel i9-13900K + RTX A6000 0.2847 0.3388 2
Intel Xeon + A100-40GB 0.2863 0.3378 64
Qwen3-8B AMD EPYC 74F3 + A100×2 0.2645 0.3083 64
AMD EPYC 9354 + RTX A6000 0.2645 0.3083 1–2
Intel i9-13900K + RTX A6000 0.2645 0.3083 2
Intel Xeon + A100-40GB 0.2673 0.3083 64
Llama-3.2-3B AMD EPYC 74F3 + A100×2 0.2436 0.2922 64
AMD EPYC 7452 + RTX 3090 0.2437 0.2922 2
Intel i9-13900K + RTX A6000 0.2445 0.2922 2
Intel Xeon + A100-40GB 0.2447 0.2907 64
Table 9: Reproducibility validation across different hardware configurations using HuggingFace backend.
Appendix D Usage Examples and Tutorials
This section provides advanced examples and tutorials for using HRET beyond the basic functionality introduced in the main paper.
D.1 Advanced Backend Usage
HRET supports multiple backend integrations beyond standard HuggingFace models:
D.1.1 LiteLLM and OpenAI-Compatible APIs
Access over 30 LLM providers and self-hosted servers:
# Using LiteLLM for cloud-based models
results = evaluator.run(
model="litellm",
model_params={
"model_name": "gpt-4",
"api_key": "your-api-key-here",
dataset="haerae_bench",
evaluation_method="string_match"
# Using OpenAI-compatible endpoints for self-hosted models
results = evaluator.run(
model="openai",
model_params={
"model_name": "local-llama-2-13b",
"api_base": "http://localhost:8000/v1",
dataset="haerae_bench",
evaluation_method="string_match"
D.2 Evaluation Methods
D.2.1 String Match and Log Probability Evaluation
HRET offers flexible text matching with various normalization options:
# String matching with normalization options
results = evaluator.run(
model="huggingface",
model_params={"model_name_or_path":"Qwen/Qwen2.5-3B-Instruct"},
dataset="haerae_bench",
evaluation_method=’string_match’,
evaluator_params={
"ignore_case": True,
"ignore_punctuation": True,
"extract_answer": True,
# Log probability evaluation for MCQA
results = evaluator.run(
model="huggingface",
dataset="haerae_bench",
evaluation_method=’log_likelihood’,
dataset_params={"base_prompt_template": "{query} ### Answer:"},
These examples demonstrate HRET’s capabilities for Korean language model evaluation beyond the basic functionality presented in the main paper. The project repository provides comprehensive tutorials and documentation in both Korean and English, including additional examples, best practices, and implementation guidelines.
Appendix E Real-World Usage
Beyond academic evaluation frameworks, HRET has been adopted in production environments across the Korean NLP industry. Authors’ affiliated organizations, including Shinhan Securities and OneLineAI, have integrated HRET into their model validation and post-training evaluation workflows, demonstrating the framework’s practical utility in real-world deployment scenarios. Additionally, the Haerae Reasoning Challenge (HRC) officially employs HRET to score all participant submissions under standardized protocols, ensuring fairness and transparency in competitive evaluations. This production adoption validates HRET’s reliability and effectiveness beyond controlled research settings, confirming its value for industry practitioners requiring robust evaluation infrastructure.
Ethics Statement
latexHRET aims to improve reproducibility and fairness in Korean LLM evaluation. However, we acknowledge several ethical considerations: (1) standardized evaluation frameworks may inadvertently favor certain model architectures or training approaches; (2) benchmark selection reflects our research perspectives and may not capture all aspects of Korean language competency; (3) language consistency enforcement, while promoting Korean language assessment, may disadvantage multilingual models in certain contexts. We encourage the community to contribute diverse benchmarks and evaluation methods to ensure broad representation of Korean language capabilities. All integrated datasets respect original licensing and guidelines.